# Explainable Machine Learning for ICU Readmission Prediction

### v1

@@ 3.2 Machine Learning Validation
The statistical analysis provided an insightful characterisation of the readmission group, contrasting it with the non-readmission group and across populations. With this information, we now move to a machine learning analysis and validation.
Table S5 shows the 47 selected features by the greedy feature selection approach, considering a balanced readmission dataset. The details of these variables representing machine learning features can be found in Table S2. Although we employed several different undersampling techniques to tackle the high level of readmission imbalance, using a balanced readmission rate presented itself as the most successful technique, which was then applied in both eICU and MIMIC IV cohorts at first. Hyper-parameter optimisation has also been applied after feature selection, aiming to improve the predictive performances of the resultant models (Table S4).
When bringing all these ML components together, our proposed pipeline achieved an area under the ROC curve (AUC) of 0.68 on multicentric eICU data under a 10-fold cross-validation procedure with a Random Forest classifier with 80 trees in a balanced readmission scenario. On the blind test (stratified 10% from eICU data), which internally validates the proposed model, consistent results were reached (AUC of 0.672) when compared to 10-fold cross-validation. When externally validated on MIMIC IV’s data, our model reached an AUC of 0.616, which demonstrates the overall generalisation capabilities of the proposed model to predict 30-day readmission. Figure 2 shows the AUC plots for these three validation schemes. The results for other performance measures are also summarised in Table S6.
@@
3.3 Calibration Analysis
Calibration metrics are presented in Table S7 for our readmission model on 10-fold cross-validation and blind testing on eICU, and external validation on MIMIC IV. Given these metrics, we present the calibration curves for these evaluation sets in Figures S4-S6.
Figure S4 depicts the calibration curve for our proposed readmission model on 10-fold cross-validation on eICU data. The calibration curve on the results in cross-validation indicates a slight overestimation of risk at predicted probabilities between 0.2 and 0.4, and a slight underestimation of risk at predicted probabilities between 0.6 and 0.7. Figure S5, in turn, shows the calibration curve for our model on a blind test set from eICU data. The calibration curve indicates a slight overestimation of risk at predicted probabilities around 0.5. Finally, Figure S6 outlines the calibration curve for our model on an external validation set, i.e., MIMIC IV. Its calibration curve indicates a slight underestimation of risk at predicted probabilities below 0.2, and an overestimation of risk at predicted probabilities above 0.4 but most significantly above 0.6. Our main conclusion is that our model is well-calibrated in all sets, demonstrating its clinical usability in an ICU setting.
@@ 3.4 Likelihood Ratio Analysis
Figure S7 summarises the likelihood ratios (LRs) for our readmission model, which performs using a classification probability threshold of 0.5. With this threshold, the model’s LRs for cross-validation on eICU, blind testing on eICU, and external validation on MIMIC IV are 1.660 (estimated change of probability of 9.628%), 1.636 (estimated change of probability of 9.348%) and 1.335 (estimated change of probability of 5.494%), respectively. These model’s results indicate its reasonable diagnostic expression.
Based on Figure S7, a threshold between 0.70 and 0.82 benefits the model the most regarding the LRs. The local and global optima based on the classification threshold 10-fold cross-validation would be 0.712 and 0.813, respectively. With 0.712, the model’s LR for 10-fold cross-validation on eICU, blind testing on eICU and external validation on MIMIC IV would be 2.905 (estimated change of probability of 20.26%), 3.326 (estimated change of probability of 22.831%) and 1.942 (estimated change of probability of 12.615%), respectively. These estimated changes show small but relevant shifts in the post-test probabilities.
With 0.813, the model reaches a maximum LR on 10-fold cross-validation on eICU (i.e., 4.256), resulting in a change of probability of readmission for 27.517%. Nevertheless, with this threshold, we have a lower LR for blind testing and external validation. In the blind test on eICU, an LR of 1.075 would be identified (with an estimated change of probability of 1.373%), while for the external validation on MIMIC IV, the model would result in an LR of 1.400 (with an estimated change of probability of readmission of 6.393%).
This analysis determines that setting properly the classification threshold is beneficial for having a better diagnostic readmission profile for ICU patients. In summary, the model’s diagnostic power depends on a proper analysis comparing and contrasting machine learning and clinical predictiveness.
@@ 3.5 Explainable Machine Learning for ICU Readmission
In addition, the SHAP tree explanation model resulted in only one hospital-related variable (see Figure 3 and Table S2), i.e., the number of minutes from unit admission time that the patient has been admitted to the hospital (i.e., the Hospital Admission Offset). In this case, there is a high correlation between lower values of Hospital Admission Offset and non-readmission, whereas relatively higher values for this variable are more related to readmission.
From ICU-related variables, the SHAP model outputs a category of unit type as important – i.e., the Unit Type Med-SurgICU, which stands for a unit type of Medical-Surgical ICU. As the variable Unit Type corresponds to a categorical variable and Med-SurgICU is one of its options, we transformed its value via one-hot encoding. In the SHAP summary plot of Figure 3, this basically means that most patients coming from Med-SurgICU are not facing readmission, while patients that are not passing through such unit type undergo readmission.
Three important demographics-related variables were also tagged in the SHAP summary from Figure 3: Age, Admission Weight and Admission Height. As seen in this figure, older people were often associated with readmission, while low values in age are noticed in non-readmitted patients. This factor is actually expected as elderly patients tend to present more comorbidities Divo et al. (2014), resulting in more readmissions. Contrarily, admission weight followed a distinct pattern, where readmitted patients had low-medium values in their readmission weight. We also noticed that readmission and non-readmission were found to be more present in high and low values of the variable admission height, respectively.
Finally, pathology-related variables, such as the Min. Partial Thromboplastin Time (First 24hs), are also at the top 20 features ranked by SHAP values. As seen in Figure 3, high Min. Partial Thromboplastin Time (First 24hs) values are more connected with readmission.
We believe these explanations, derived by SHAP for the proposed readmission model, regarding the top 20 most impactful variables might serve as a new source of guidelines while assessing patients through ICU, especially while discharging patients. Although clinical practice and experience will always be compulsory in such conditions, we believe machine learning models might provide proper support for them.
@@ 4.1 Key Findings
Although it is challenging to map and match common variables or characteristics across different cohorts, our study drove important clinical insights while analysing eICU and comparing it to MIMIC IV (Section 3.1). We discovered how heterogeneous the readmission monocentric and multicentric populations are in general. For example, the readmitted patients in MIMIC IV were treated more frequently with ventilation and vasopressors, although renal replacement therapies were more found in patients readmitted to eICU. Such contrasts in the cohort populations highlight the ability of our standardised pipeline to deal with heterogeneous readmission data.
@@
Our proposed machine learning pipeline learned proper patterns on multicentric ICU data, consequently generalising well on an independent blind test set over the same data (Section 3.2). Our proposed machine learning model captured the essence of readmission, being able to transfer similar predictive performance to external validation on monocentric ICU data (Section 3.2). Overall, this shows learning on data coming from multiple ICUs has its limitations but may lead to generalisable predictive performance if done adequately.
@@
A reasonable calibration level is also an essential aspect of the proposed readmission model (Section 3.3). As a result, our proposed model shares meaningful clinical decision-making due to its ability to calculate individualised probabilistic estimates of readmission Munn et al. (2022). For example, clinicians may decide that patients with a risk of readmission at least double that of the average risk in their ICU population should be flagged for routine monitoring following discharge from ICU. If the patient is moved to the ward, this may take the form of a brief clinical review by an intensivist each day following transfer to identify early signs of clinical deterioration. This could potentially facilitate the escalation of ward-based measures to optimise the patient’s management and avoid ICU readmission, or it could facilitate better preparation and planning such that unplanned readmissions are minimised.
@@
A good diagnostic level, when we increase the classification probability threshold, complements the model’s predictive performance (Section 3.4). At the threshold(s) of 0.712 (and 0.813), our model returns the most reliable likelihood ratio among cross-validation, blind testing and external validation, yielding trustworthy readmission predictions.
Finally, we highlight how the 20 most impactful features are linked with readmission prediction through their explainability (Section 3.5). These explanations might be used to support the clinical team. In fact, some of the insights provided by the interpretation of the readmission model’s explanation are corroborated by the literature. We have highlighted these analyses and observations for some of the important variables displayed by SHAP values (Figure 3) next:
@@
    Low serum albumin levels in ICU patients have already been linked with death, length of stay, and readmission Rojas et al. (2018); Herrmann et al. (1992). These results from previous work are congruent with our analysis based on Figure 3.
@@
    High values of blood urea nitrogen (BUN) have also been associated with readmission and mortality in past studies Rojas et al. (2018); Arihan et al. (2018); Li et al. (2021), representing an important feature to characterise critically ill patients admitted to ICU.
@@
    Previous studies also mapped low hemoglobin levels preceding ICU discharge, which may provide important insights for avoiding readmission Hammer et al. (2021); Warner et al. (2022); Kerfeld et al. (2020).
@@
    Elderly patients usually are observed to be more associated with ICU readmission Lee et al. (2022); Glans et al. (2020); Pedersen et al. (2017).
  • 5.
    The patients’ weight is strongly correlated to readmission. While gaining weight through an ICU readmission is linked to non-readmission in patients with heart failure Howie-Esquivel et al. (2019), weight and weight change are common factors associated with mortality, duration of mechanical ventilation, and length of stay in the ICU You et al. (2013); Mishra et al. (2021). Therefore, looking at the patient’s admission weight and tracking their respective weight during ICU admission is a relevant aspect before discharging a patient from the ICU.
Overall, this shows our proposed model was able to learn and capture good insights on eICU data, being compatible with results and discussion in the ICU readmission literature.
4.2 Strengths and Limitations
One of the limitations of this work is the lack of comparable variables between multicentric eICU and monocentric MIMIC IV, including complete APACHE IV scores, which are not present in MIMIC IV. Nevertheless, we matched 168 common variables between the two cohorts, constituting both patient- and institutional-level information, building a key strength of our characterisation.
Another aspect is that several variables presented a large portion of missing data (more than 80%), bringing an extra level of difficulty to perform our methodology. We handled missing data by excluding columns with more than 20% of missing values and performing imputation in the other cases. This decision was taken because we considered we would include a high proportion of bias and noise if we impute columns with a great frequency of missingness. Although missingness might be informative Tan et al. (2023), the generalisation reached by the proposed predictive readmission model showed the robustness of our ensemble-based methodology to deal with different noise levels – from both missingness and utilised imputation approach. We plan in future work to better understand missingness across eICU and MIMIC IV as a way to improve machine learning modelling.
In addition, we observed the challenge of finding comparable metrics for the severity of the illness to describe the population between eICU and MIMIC. Therefore, our attempt resumed using the variables included in most severity of illness scores in our data description. Given data restrictions, we believe this was the best way to compare the severity of illness between both cohorts.
Finally, we noticed an overall heterogeneity of the studied populations, including imbalance due to rare occurrences of readmission, where the eICU cohort presents 4.04% of readmission cases while MIMIC IV has more than double (8.74%). We successfully dealt with the imbalance by employing appropriate performance metrics on top of the feature selection method Chicco and Jurman (2020). Our choice for the set of metrics considered the four confusion matrix categories (true positives, false negatives, true negatives, and false positives), setting proportional weights based on the size of the readmission group and the size of the non-readmission group in the datasets. Undersampling techniques also assisted with diminishing the imbalance hardness in our study.
4.3 Relationship to Previous Studies
Multiple predictive models were developed for readmission with different AUC performances (see Table S1). Nevertheless, we believe that most of these models lack clinical applicability. The main reason behind this is that those related works prefer highlighting their good predictive performances rather than understanding readmission well and bridging meaningful clinical and machine learning knowledge.
Although performing similarly to the literature, most of the works are focused on monocentric data, except for Badawi and Breslow Badawi and Breslow (2012) and Mohanty et al. Mohanty et al. (2022). However, both works lack clinical decision-making and are not well-validated with internal and external data sets. Apart from Mohanty et al. employing explainable machine learning in their proposed models, providing insights for their results, they remain local to the Epic EHR database employed by them.
Hegselmann et al.’s work Hegselmann et al. (2022) is the closest to our method. Their readmission model is built from University Hospital Münster (UKM) data and externally tested on MIMIC IV. Even though it presents a certain level of similarity to our work, their work is set in a controlled scenario, where UKM data is employed to build the model. Furthermore, both UKM and MIMIC IV are monocentric, presenting, therefore, a lower level of difficulty than multicentric eICU. In our method, we learn a readmission model from multicentric data to generalise on monocentric data. We show our method works well in translating from one to another.
Funding
Conflict of Interest
None declared.
Contributions
Alex G. C. de Sá: Conceptualisation, Draft figure proposals, Literature review, Methodology, Machine Learning - Experimental planning and execution, and results analysis, Visualisation, Writing - Original draft preparation and supplementary materials, Reviewing, Editing, and Validation (machine learning perspective);
Daniel Gould: Conceptualisation, Literature review, Visualisation, Investigation, Writing - introduction and results, Analysis, Reviewing, Editing, and Validation (clinical and machine learning perspectives);
Anna Fedyukova: Data collection, Machine Learning - Experimental planning and execution, and results analysis, Visualisation, Writing- Methodology and results, supplementary materials, and Validation (machine learning perspective);
Mitchell Nicholas: Literature review, Reviewing, Editing, and Validation (clinical perspective);
Calvin Fletcher: Writing - introduction, Validation (clinical perspective), Reviewing, and Editing;
Lucy Dockrell: Writing, Validation (clinical perspective), Reviewing, and Editing;
David Pilcher: Supervision, Writing, Validation (clinical perspective), Reviewing, and Editing;
Daniel Capurro: Conceptualisation, Supervision, Writing, Reviewing and Editing;
David B. Ascher: Supervision, Writing, Reviewing, Editing and Validation (clinical and machine learning perspectives);
Khaled El-Khawas: Conceptualisation, Supervision, Supervision, Writing- Methodology and Results, and Validation (clinical perspective);
Douglas E. V. Pires: Conceptualisation, Supervision, Writing, Reviewing, Editing, and Validation (clinical and machine learning perspectives);
@@ 4.3 Relationship to Previous Studies
Supplementary Materials are available at https://tinyurl.com/36f333sc.

### v4

@@ 3.2 Machine Learning Validation
The statistical analysis provided an insightful characterisation of the readmission group, contrasting it with the non-readmission group and across populations. With this information, we can now move to a machine learning analysis and validation.
Table S5 shows the 47 selected features by the greedy feature selection approach using a Random Forest classifier and considering a balanced readmission dataset. The details of these variables representing machine learning features can be found in Table S2. Although we employed several different undersampling techniques to tackle the high level of readmission imbalance, using a balanced readmission rate presented itself as the most successful technique, which was then applied in both eICU and MIMIC IV cohorts at first. Hyper-parameter optimisation was also applied after feature selection, with the aim of improving the predictive performances of the resultant models (Table S4).
When bringing all these ML components together, our proposed pipeline achieved an area under the ROC curve (AUC) of 0.68 on multicentric eICU data under a 10-fold cross-validation procedure with a Random Forest classifier with 80 trees in a balanced readmission scenario. It is worth noting that this classification model was considered the best against 9680 other machine learning models in terms of a trade-off among AUC, MCC and recall metrics on 10-fold cross-validation. Explainability was also a key point in selecting the model. Therefore, the selected variables (features) were taken into consideration while selecting the predictive model itself. On the blind test (stratified 10% from eICU data), which internally validates the proposed model, consistent results were reached by the predictive model (AUC of 0.672) when compared to 10-fold cross-validation. When externally validated on MIMIC IV’s data, our model reached an AUC of 0.616, which demonstrates the overall generalisation capabilities of the proposed model to predict 30-day readmission. Figure 2 shows the AUC plots for these three validation schemes. The results for other performance measures are also summarised in Table S6.
@@
3.3 Calibration and Likelihood Analysis
Based on the performed Calibration Analysis (Supplementary Results, Table S7 and Figures S4-S6), the calibration curves, defined from our 30-day ICU readmission model during cross-validation, blind testing and external testing, indicated slight overestimation/underestimation. Therefore, our main conclusion is that our model is well-calibrated in all sets, demonstrating its clinical usability in an ICU setting.
@@ 3.4 Likelihood Ratio Analysis
In addition, our 30-day ICU readmission model faced an analysis based on Likelihood Ratios (see Supplementary Results and Figure S7), which highlighted the importance of properly setting the classification probability threshold. This analysis step is beneficial for having a better diagnostic readmission profile for ICU patients. In summary, the ICU predictive model’s diagnostic power depends on a proper analysis of these thresholds, which was properly defined after the Likelihood Ratio Analysis (Supplementary Results).
@@ 4 Discussion
Although it is challenging to map and match common variables or characteristics across different cohorts, our study drove important clinical insights while analysing eICU and comparing it to MIMIC IV (Section 3.1). We discovered how heterogeneous the readmission monocentric and multicentric populations are in general. For example, the readmitted patients in MIMIC IV were treated more frequently with ventilation and vasopressors, although renal replacement therapies were more common in patients readmitted to eICU. Such contrasts in the cohort populations highlight the ability of our standardised pipeline to deal with heterogeneous readmission data.
@@
Our proposed machine learning pipeline learned proper patterns on multicentric ICU data, consequently generalising well on an independent blind test set over the same data (Section 3.2). Our proposed machine learning model captured the essence of readmission, being able to transfer similar predictive performance to external validation on monocentric ICU data (Section 3.2). Overall, this shows that learning from data coming from multiple ICUs has its limitations but may lead to generalisable predictive performance if done adequately.
@@
A reasonable calibration level is also an essential aspect of the proposed readmission model (Supplementary Results). As a result, our proposed model shares meaningful clinical decision-making due to its ability to calculate individualised probabilistic estimates of readmission [36]. For example, clinicians may decide that patients with a risk of readmission at least double that of the average risk in their ICU population should be flagged for routine monitoring following discharge from ICU. If the patient is moved to the ward, this may take the form of a brief clinical review by an intensivist each day following transfer to identify early signs of clinical deterioration. This could potentially facilitate the escalation of ward-based measures to optimise the patient’s management and avoid ICU readmission, or it could facilitate better preparation and planning such that unplanned readmissions are minimised.
@@
A good diagnostic level, when we increase the classification probability threshold, complements the model’s predictive performance (Supplementary Results). At the threshold(s) of 0.712 (and 0.813), our model returns the most reliable likelihood ratio among cross-validation, blind testing and external validation, yielding trustworthy readmission predictions.
Finally, we highlight how the 20 most impactful features are linked with readmission prediction through their explainability (Section 3.3 and Supplementary Results) [17, 37, 38, 39, 21, 40, 41, 42, 43, 44, 45, 46, 47]. Overall, this shows our proposed model was able to learn and capture good insights on eICU data, being compatible with results and discussion in the ICU readmission literature. Supplementary Discussion provides more details on the model’s explainability, strength and limitations of this work and its relationships to other studies.
Authors’ contributions
Alex G. C. de Sá: Conceptualisation; Formal analysis; Investigation; Methodology; Project administration; Resources; Software; Validation; Visualisation; Writing – original draft; Writing – review & editing.
Daniel Gould: Conceptualisation; Formal analysis; Investigation; Methodology; Project administration; Validation; Visualisation; Writing – original draft; Writing – review & editing.
Anna Fedyukova: Data curation; Formal analysis; Methodology; Software; Validation Visualisation; Writing – original draft.
Mitchell Nicholas: Validation; Writing – review & editing.
Calvin Fletcher: Validation; Writing – original draft; Writing – review & editing.
Lucy Dockrell: Validation; Writing – review & editing.
David Pilcher: Supervision; Validation; Writing – review & editing.
Daniel Capurro: Conceptualisation; Validation; Writing – review & editing.
David B. Ascher: Supervision; Validation; Writing – review & editing.
Khaled El-Khawas: Conceptualisation; Investigation; Methodology; Project administration; Supervision; Validation; Writing – original draft; Writing – review & editing.
Douglas E. V. Pires: Conceptualisation; Methodology; Project administration; Supervision Validation; Writing – original draft; Writing – review & editing.
Acknowledgements
Statement on Conflicts of Interest
None declared.
Summary Table
@@ Acknowledgements
    We propose a standardised and explainable machine learning pipeline for 30-day readmission prediction in multicentric eICU and monocentric MIMIC IV cohorts.
@@
    Our random forest-based readmission model yields overall good calibration and generalisable predictive performance on internal and external validation sets.
@@
    Calibration and likelihood ratio analysis also demonstrated the proper clinical (diagnostic) usability of our proposed model, even with the challenging nature of the studied readmission datasets.
@@
    The derived explanations from the proposed readmission model provide invaluable guidelines, which might be helpful for ICU clinicians’ decision-making while discharging patients.
@@
Supplementary Materials are available at https://tinyurl.com/fne6xv2t.